40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15172     Entry time: Wed Jan 29 00:29:43 2020
Author: gautam 
Type: Update 
Category: LSC 
Subject: locking 2020 

The goal tonight was to go through the locking scripts to see if I could recover the state from November 2019, when I could have the arm lengths controlled by ALS, and sit at zero CARM offset with the PRMI remaining locked and the arm powers fluctuating between 0-300. The IR-->ALS transitions went smoothly tonight, and the PRMI locking was also fairly robust when the CARM offset was large, but was less good when reduced to 0. Nevertheless, it is good to know that the system can be restored to the state it was late last year. Next step is to figure out how to keep the PRMI locked and get the AO path engaged, this was what I was struggling with before the new year.

Attachment 1: PRFPMI_2020Jan.png  172 kB  | Hide | Hide all
PRFPMI_2020Jan.png
ELOG V3.1.3-